Let X1 and X2 denote the number of three point attempts required respectively by Jack and Jill for to get their first successful three pointer. When Jack joined Jill in a three point shoot-out, Jill had already made k unsuccessful attempt. Assume that Jack and Jill both are equally good in their ability to hit three point shot and that the probability of successful three point shot for each of them is 0.3. What is the probability that Jill will not have made a successful three point shot in another s attempts after jack joined him. What is the probability that Jack will not have made successful three point shot in s attempts after he joined the Jill.
Answer #1

Answer:

a)

Given,

Probability of successful three point shot = 0.3

i.e.,

P(Success) = 0.3

Here Jill had already made k unsuccessful attempt

Required probability = ( 1 - 0.3)^s

= 0.7^s

Hence the probability that Jill will not have made a successful three point shot in another s attempts after jack joined him is 0.7^s

Here it doesn't matter whether there are k unsuccessful attempts as Jack hadn't joined

b)

To give the probability that Jack will not have made successful three point shot in s attempts after he joined the Jill

i.e.,

= 0.7^s